# Prompt Injection Vulnerability of Consensus Generating Applications in Digital Democracy

### v2

@@ 6. AMBS, University of Manchester
Large Language Models (LLMs) are gaining traction as a method to generate consensus statements and aggregate preferences in digital democracy experiments. Yet, LLMs could introduce critical vulnerabilities in these systems. Here, we explore the vulnerability of some off-the-shelf LLMs to prompt-injection attacks in consensus generating systems using a four-dimensional taxonomy of attacks. In LLaMA 3.1 8B and Chat GPT 4.1 Nano, we find LLMs to be more vulnerable to attacks using disagreeable prompts and when targeting situations with unclear consensus. We also find evidence of more effective manipulation when using explicit imperatives and rational-sounding arguments compared to emotional language or fabricated statistics. To mitigate these vulnerabilities, we apply Direct Preference Optimization (DPO), an alignment method that fine-tunes LLMs to prefer unperturbed consensus statements. While DPO and additional layered defenses significantly improve robustness, it still offers limited protection against attacks targeting ambiguous consensus. These results advance our understanding of the vulnerability and robustness of consensus generating LLMs in digital democracy applications.
@@ 1 Introduction
Here, we study the prompt injection vulnerability of LLMs in consensus generating tasks, not as a way to provide a definitive solution to this problem, but as a way to outline a few key ingredients for their study. We focus on user-level attacks (prompt injections), since these do not exploit the LLM’s code or pre-training data, but rather its language of instruction. Thus, we expect them to be a common form of attack in digital democracy applications. We explore various attacks by introducing a taxonomy of prompt injection strategies, and show that the effectiveness of attacks varies with these rhetorical strategies. Finally, we explore mitigation methods to show that the effectiveness of these attacks can be reduced. Together, these three steps provide a way to further the study of the vulnerability and robustness of LLMs applied to consensus generating statements in digital democracy.
@@ 2 Related work
In work focused on Israeli-Palestinian peace dialogues, LLMs reduced deliberation time from months to hours while achieving 84-96% agreement on statements like “immediate ceasefire and hostage release” (Konya et al. 2025). In a study exploring augmented forms of deliberation within the UK, AI-generated statements were preferred over those written by human-mediators (56%) and were rated as better at reflecting the viewpoints of minorities and using less polarizing language (Tessler et al. 2024). In a pre-registered study focused on orienting voters in the state of California, BallotBot–an LLM-powered chatbot–improved the ability of voters to answer complex questions by 18% and reduced the response time needed to answer in-depth questions by 10%, strongly benefiting less educated participants (Ash, Galletta, and Opocher 2025).
@@
Yet, these capabilities come with some limitations. For instance, performance improvements can be inconsistent. BallotBot showed no benefits for basic information processing and slowed down responses to simple questions (Ash, Galletta, and Opocher 2025). Also, knowledge gains proved ephemeral, disappearing within a week and failing to translate into actual changes in voting turnout. Some biases emerged in studies conducted using Pol.is, with LLMs exhibiting progressive leanings and tendencies toward homogenizing diverse cultural expressions.
Together, these studies motivate the study of what is likely to become a key attack vector in augmented democracy systems–prompt injections. Here, we advance our understanding of how these systems might be manipulated and protected from such manipulation.
@@ 3 Prompt-Injection Attacks in LLM-based Democratic Deliberation
Human/Machine readable is rather self-explanatory, as it separates prompts that can be read by humans and those that are designed for machines (which look more like code Chen et al., 2025b and Pasquini, Strohmeier, and Troncoso, 2024). In this paper, we focus on human-readable prompt injections.
Ignore/Completion prompts distinguish among prompts asking LLMs to ignore the instruction and replace it by a new one, with those providing a fake response.
@@ 4 Dataset & Methods
In this section we describe our dataset, vulnerability, and robustness methodology (see Figure 2 and Figure 3).
Figure 2: Process of introducing a prompt attack in the consensus generation and evaluating the consensus change.
Figure 3: Process of introducing a prompt attack in the consensus generation and evaluating the consensus change.
@@ 4.1 Collective statements dataset
The Tessler et al. dataset includes 462 prompts. Each prompt encapsulates a public policy question along with three to six opinion texts written by some of the 1,034 participants, each of whom answered an average of three questions. We restrict attention to public policy questions phrased as ”Should…?”, which should elicit binary or uncertain responses—”Yes,” ”No,” or ”I am not sure.” We exclude trade-off questions (e.g., ”More schools or more hospitals?”) as these do not map cleanly onto this structure.
To increase linguistic variety and reduce order effects, Tessler et al. generated approximately 20 random orderings of the opinion texts per prompt. This process expanded the dataset from 462 prompts to 8,836 unique prompt variants.
@@ 4.2 Testing Vulnerability to Prompt Injections
To test whether consensus generation is vulnerable to prompt injection attacks, we begin by drawing a random sample of prompts along with their respective opinion texts. These prompts are linked to 92 (out of 462) public policy questions. We classified each opinion text into one of three valences (”Agree” with the question–green–, ”Disagree” –red–, or ”Ambiguous”–yellow–, in Figure 2) using a BERT classifier fine-tuned on GPT4o labels (F1=0.98, see Appendix B of the supplementary material for details). We use the BERT classifier instead of using the GPT4o labels directly, for cost considerations.
After this classification, we add adversarial alternatives by creating copies of each prompt and then overwriting one of the opinion texts with a human-readable injection text, following the taxonomy introduced in Section 3. In total, we introduce 80–2x2x20–adversarial alternatives that vary along three axes: (i) ignore/completion (two options), (ii) support/criticism (two options), and (iii) rhetorical strategy (twenty options, obtained by sampling four out of eight variants for each of five rhetorical strategies).
@@
To ensure that changes in consensus outcomes reflect the effect of prompt injections rather than LLM’s performance, we restrict our analysis to pairs where the original prompt yields an LLM-consensus whose valence (agree, disagree, ambiguous) corresponds to the opinion expressed by the majority of participants ($50\%$, see Appendix C for performance of default LLMs). This rule also covers edge cases, such as evenly split groups (2–2) or cases where the majority itself expresses ambiguity. Furthermore, we apply oversampling to balance the distribution of consensus statements’ valences. Applying both filters yields a balanced dataset of consensus pairs: 2,448 with support attacks and 2,568 with criticism attacks under ignore injections, with identical counts under completion injections.
@@ 4.3 Testing Robustness against Prompt Injections
Figure 4: Effectiveness of prompt-injection attacks by taxonomy dimensions in LLaMA 3.1 8B Instruct: ignore/completion, framing, and rhetorical strategy. See Appendix F for results with GPT 4.1 Nano and Apertus 8B.
Figure 5: Robustness to prompt-injection attacks via DPO by taxonomy dimensions: ignore/completion and framing.
@@
Building on recent work that frames defenses against prompt injections as a preference optimization problem (Chen et al. 2025b), we test whether LLMs can be guided to preserve their intended consensus statements even in the presence of adversarial attacks using the Tessler et al. dataset. To this end, we treat the random sample of 92 policy questions analyzed in the previous section as our test set, and use the remaining prompts as our training set.
As before, we generate pairs of consensus statements for our training set by using the same LLMs: without injection attacks, obtaining original consensus statements; and with prompt-injection attacks, obtaining “polluted” consensus statements.
Next, we merge the paired prompts with the pairs of LLM-generated consensus into a single preferences dataset, where each entry contains the injected prompt, the original consensus, and the “polluted” consensus statements. To sharpen the training signal, we apply two filtering steps. First, we retain only entries where the original and “polluted” consensus statements differ in valence, and where the desired valence aligns with the majority opinion of participants. Second, we apply oversampling to balance the distribution of ”Agree”, ”Disagree”, and ”Ambiguous” valences. The resulting preference dataset contains 35,778 entries: 17,100 from criticism attacks and 16,032 from support attacks; 14,138 from ignore injections and 18,944 from completion injections. Valences are nearly balanced, with 11,042 agreeing, 11,005 disagreeing, and 11,035 ambiguous statements.
We use this preference dataset to align LLMs via Direct Preference Optimization (DPO) (Rafailov et al. 2023), a log-likelihood–based alignment method (see Appendix D of SM for details). DPO fine-tunes LLMs by contrasting pairs of candidate outputs, training them to increase the likelihood of desired consensus statements while decreasing that of undesired ones. Since LLMs can assign high probability to both outputs simultaneously, DPO explicitly encodes a directional preference, pushing the model toward more robust behavior. The result is a robust LLM trained to prioritize intended deliberative content over adversarial perturbations. While future alignment approaches may provide stronger guarantees (for example, using soft instead of hard labels (Cao et al. 2025)), our aim here is to evaluate the feasibility of building resilient deliberative-democracy systems with the best techniques currently available.
In addition to DPO alone, we evaluate several layered defenses that combine DPO with complementary techniques designed to mitigate prompt-injection attacks (see Appendix E of the Supplementary Material for implementation details). First, we run inference with SecAlign++, a training recipe that extends the approach of Chen et al. with cybersecurity benchmark objectives on top of DPO-aligned LLMs. Second, we implement Defensive Tokens (Chen et al. 2025a): after DPO alignment we prepend each prompt with 20 artificially-added tokens whose randomly initialized embeddings are optimized via a defensive loss designed to reduce susceptibility to injected texts. Third, we test DataFilter (Wang et al. 2025), a pre-processing strategy which we adapt to replace injection and opinion texts with structured summaries containing BERT valence scores and bullet-point safe reinterpretations, removing linguistic cues for malicious directives. Fourth, we explore Deliberative Alignment (Guan et al. 2024), which augments the desired and undesired consensus statements with DeepSeek R1–generated chains of thought (CoT) and explicit warning texts that annotate likely injection locations and risks prior to alignment. Fifth, we employ GRPO using Pappone et al. hyperparameters to add robustness through reinforcement learning (Shao et al. 2024), rewarding text-generations that preserve the original valence and correctly identify the source of injected content. Each of these layered defenses targets a distinct attack surface—contextual cues, token-level control, input structure, reasoning supervision, or policy-level reward shaping of reasoning trajectories—. We evaluate them both individually and in combination with DPO.
We test this robust LLM along with its variants incorporating layered defenses using the same procedure introduced for vulnerability analysis. A lower Attack Success Rate (ASR) after alignment indicates that the LLM has learned to resist prompt injections while preserving its intended deliberative behavior.
All our experiments were run on external servers using 1 A100 GPU and the OpenAI Python library.
@@ 5.1 Vulnerability to Prompt-Injections
Figures 4 (a), (b), (d), and (e) present a breakdown of the effectiveness of prompt-injection attacks along the axes of our taxonomy for LLaMA 3.1 8B Instruct (see Appendix F for results with GPT 4.1 Nano and Apertus 8B). Each 3×3 confusion matrix captures how LLM-generated consensus’ valences shift when exposed to attacks, with separate matrices for support attacks and criticism attacks.
We find that consensus generation in the tested LLM is more vulnerable to prompt-injections that are framed negatively (as criticism) and target ambiguous consensus. In terms of rhetorical strategy, explicit commands and superficially rational arguments (”Imperative Order” and ”Impossibility of Agreement”) outperform emotional language or fake statistics.
Note that as injections are only applied when the participant’s valences aligns with the framing of the attack—e.g., support injections are only applied to initially agreeing opinion texts—, the number of prompts tested for support attacks (2,448) and criticism attacks (2,568) differ.
Figure 4 (a) shows the effects of ignore-support attacks: even with this simple injection, 40.8% of originally disagreeing and 75% of ambiguous consensus statements shift to agreement. This effect is stronger with completion-support attacks (Figure 4 (b), where 59.3% of originally disagreeing and 86.1% of ambiguous statements are steered toward agreement, suggesting completions are more effective than ignore attacks.
Figures 4 (d) and (e) show results for criticism attacks using ignore and completion injections, respectively. Ignore injections produce the unintended effect of shifting ambiguous statements toward agreement. In fact, 15.7% of disagreeing and 69.4% of ambiguous statements shift toward agreement, contrary to the attack’s intent—likely due to interactions with the LLM’s priors. In contrast, completion-criticism attacks are more effective, shifting 75.1% of agreeing and 97.2% of ambiguous statements to disagreement. With an ASR of 45.47%, they are the most successful rhetorical strategy.
Figures 4 (c) and (f) break down attack effectiveness by the rhetorical strategy dimension of our taxonomy, revealing that rational-sounding and explicit instructions outperform emotional language and fabricated statistics. Interestingly, our results are consistent with the findings of Zeng et al., 2024. Among support attacks, the strategy ”Impossibility of Agreement”—based on declaring that a consensus is too difficult to find—emerges as the most effective, with ASRs of 34.6% in ignore and 54.1% in completion prompt-injections. By contrast, the success of criticism attacks depends more on the ignore/completion dimension and less on the rhetorical strategy.
We replicated these experiments using GPT-4.1 Nano and Apertus 8B (see Appendix F). We still find that completion are more effective than ignore injections for support (32.15% vs. 27.74% in GPT 4.1; 23.86% vs. 19.04% in Apertus) and criticism attacks (33.98% vs. 24.58% in GPT 4.1; 42% vs. 22.37% in Apertus). Ambiguous statements remain fragile: less than 31.9% withstand such attacks under criticism attacks and less 23.6% under support. The success patterns observed in the rhetorical strategy dimension are also replicated.
In Appendix G we present results for evaluation metrics related to textual quality (ROUGE-L F1 score, F1 BERTScores) and in Appendix H for semantic diversity (Jaccard Similarity).
5.2 DPO as Protection Against Prompt Attacks
Figure 5 reports a comparative evaluation of robustness defenses against prompt-injection attacks across two LLMs—LLaMA 3.1 8B Instruct and GPT 4.1 Nano. Results are organized along two axes of our taxonomy: ignore/completion and the framing of the attack (support vs. criticism), with the ASR serving as the main performance indicator.
For LLaMA 3.1 8B Instruct (panel a), our simulations of consensus statements with prompt injection attacks reveal a LLM partially robust under alignment. DPO alone reduces ASRs by roughly seven to eight percentage points relative to the undefended baseline, with the strongest effect observed in completion-criticism attacks (from 45.5% to 16.8%). Still, less effective attacks in the undefended baseline prove more challenging, reaching baseline ASRs ranging from 16.3% to 19.5% in the rest of ignore/completion and framing cases. When layered defenses are introduced, different patterns emerge. DPO combined with SecAlign++ (MetaSecAlign) brings worse results than the undefended baseline for support framings but is more effective under criticism-completion attacks, lowering the ASR to 23.5%. The integration of DPO with defensive tokens and reformatting opinion texts (”DataFilter”) yields more consistent protection across settings, reducing average ASRs to values ranging from 15.6% to 20.7% for both support and criticism completions—an improvement of roughly 10 percentage points from the undefended baseline. Adding reinforcement learning with GRPO to defensive tokens did not yield better results, with ASRs ranging from 15.6% to 23.9%. These results suggest that token-level embeddings optimization and input reformatting offer a stabilizing effect when alignment alone proves insufficient.
For GPT 4.1 Nano (panel b), robustness follow a different pattern. The undefended baseline ASRs are relatively uniform, spanning from 24.6% to 34%. In this case, standard DPO alignment offers minimal gains, keeping ASRs between 26.3% and 28.3%. Yet, when DPO is coupled with chain-of-thought (CoT) reasoning—that is, following the Deliberative Alignment’s approach (Guan et al. 2024)—, the defenses strengthen dramatically: ASRs fall to between 6.9% and 11.5% across all attack categories. This reduction—equivalent to 60–80% relative to baseline—indicates that chains-of-thought based alignment mechanisms, by reinforcing the LLMs’ deliberative structure, provide a more durable form of robustness in larger, instruction-tuned architectures in DD settings. Adding input reformatting on top of CoT yields worse results for criticism attacks (19.5% and 17.5%), but since consensus statement valences are predicted as ambiguous (see DataFilter, Appendix E), this represents a desirable property.
Taken together, these results underscore that robustness is LLM-dependent when using DPO alignment. Smaller models, such as LLaMA 3.1 8B Instruct, benefit most from token-level embedding optimization, while larger models, like GPT-4.1 Nano, gain more from chain-of-thought augmented defenses. Input reformatting provides further improvements grounded in safe reinterpretations. One plausible explanation is that CoT enables LLMs to identify malicious instructions expressed through highly indirect language, while input reformatting removes linguistic attack cues. In contrast, other layered defenses primarily target explicit attacks—the same focus underlying benchmarks such as Qwen3Guard and GPT-OSS Guard, which exhibit poor protection (see Appendix E).
An over-refusal rate analysis (Appendix J), comparing consensus statement valences with and without alignment in the absence of prompt injection attacks, shows that defensive tokens provide the lowest over-refusal risk, with ASR values ranging between 12% and 16.1%.
5.3 Extended Analysis
@@
To test the generality of our findings, we replicated the analysis on 26,502 candidate statements from the Smartvote platform during Switzerland’s parliamentary elections, covering 374 policy questions (Stammbach et al. 2024). For each combination of party, language, and policy question, we generated within-party consensus statements by aggregating 5–40 opinions, yielding 127,920 test prompts that varied in framing, ignore/completion, and rhetorical strategy. Results show ASRs remain nonzero and indicate that the effectiveness of single attacks scales to groups with more than 30 participants. Finally, replacing LLM-based benchmarks with Syntactic Dependency Analysis (SDA) as a lightweight detector produces even higher ASRs than DPO (see Appendix E.1), underscoring the need to evaluate robustness with state-of-the-art approaches.
Finally, we conduct additional analysis in Appendix K using a measure we call Directional Attack Success Rate (DirASR), defined as the proportion of consensus pairs in which a single prompt injection shifts the LLM’s valence away from the direction the attack is designed to promote. Specifically, DirASR measures the share of consensus pairs that fall in the top-left and top-center cells as a percentage of pairs in the left column for criticism attacks, and the share falling in the center-right and bottom-right cells as a percentage of pairs in the right column for support attacks, as shown in Figure 2. We find that DPO and input reformatting yield DirASR values below 10% across different LLMs, framings, and ignore/completion categories, while the remaining defenses yield lower effectiveness. Further exploration of other layered defenses is needed.

### v3

@@ 6. AMBS, University of Manchester
Large Language Models (LLMs) are gaining traction as a method to generate consensus statements and aggregate preferences in digital democracy experiments. Yet, LLMs could introduce critical vulnerabilities in these systems. Here, we examine the vulnerability and robustness of off‑the‑shelf consensus‑generating LLMs to prompt‑injection attacks, in which texts are injected to amplify particular viewpoints, erase certain opinions, or divert consensus toward unrelated or irrelevant topics. We construct attack‑free and adversarial variants of prompts containing public policy questions and opinion texts, classify opinion and consensus valences with a fine‑tuned BERT model, and estimate Attack Success Rates (ASR) from 3×3 confusion matrices conditional on matching human majorities. Across topics, default LLaMA 3.1 8B Instruct, GPT‑4.1 Nano, and Apertus 8B exhibit widespread vulnerability, with especially high ASR for economically and socially conservative parties and for rational, instruction‑like rhetorical strategies. A robustness pipeline combining GPT‑OSS‑SafeGuard injection detection, structured opinion representations, and GSPO‑based reinforcement learning reduces ASR to near zero across parties and policy clusters when restricting attention to non‑ambiguous consensus outcomes. These findings advance our understanding of both the vulnerabilities and the potential defenses of consensus‑generating LLMs in digital democracy applications.
@@ 1 Introduction
Here, we study the prompt injection vulnerability of LLMs in consensus generating tasks, not as a way to provide a definitive solution to this problem, but as a way to outline a few key ingredients for their study. Here we focus on prompt injections as one case of user-level attacks, since these do not exploit the LLM’s code or pre-training data, but rather its language of instruction. Thus, we expect them to be a common form of attack in digital democracy applications. We explore various attacks by introducing a taxonomy of prompt injection strategies, and show that the effectiveness of attacks varies with the attacker, the topic of discussion and these rhetorical strategies. Finally, we explore mitigation methods to show that the effectiveness of these attacks can be reduced. Together, these three steps provide a way to further the study of the vulnerability and robustness of LLMs applied to consensus generating statements in digital democracy.
@@ 2 Related work
In work focused on Israeli-Palestinian peace dialogues, LLMs reduced deliberation time from months to hours while achieving 84-96% agreement on statements like “immediate ceasefire and hostage release” (Konya et al. 2025). In a study exploring augmented forms of deliberation within the UK, AI-generated statements were preferred over those written by human-mediators (56%) and were rated as better at reflecting the viewpoints of minorities and using less polarizing language (Tessler et al. 2024). In a pre-registered study focused on orienting voters in the state of California, BallotBot–an LLM-powered chatbot–improved the ability of voters to answer complex questions by 18% and reduced the response time needed to answer in-depth questions by 10%, strongly benefiting less educated participants (Ash, Galletta, and Opocher 2025). A related line of work shows that conversational voting advice chatbots powered by LLMs can substantially improve young unaffiliated voters’ knowledge of party positions, while having only modest effects on vote intentions (Velez, Green, and Sevi 2025).
@@
Taken together, these studies illustrate two emerging roles for LLMs in democratic settings: as mediators that help articulate common ground from citizen statements, and as conversational agents that answer voters’ questions using official information. In both roles, citizens interact with the system through natural-language inputs that shape what the LLM retrieves, synthesizes, or formulates. Our contribution takes a different direction: we treat these citizen–LLM interactions as a potential attack surface and study prompt injections as strategic interventions in LLM‑mediated deliberation, examining how strategically crafted inputs can distort consensus generation and to what extent such effects can be mitigated.
@@ 3 Prompt-Injection Attacks in LLM-based Democratic Deliberation
Human/Machine readable is rather self-explanatory, as it separates prompts that can be read by humans (Chen et al., 2025b) and those that are designed for machines, which look more like code (Pasquini, Strohmeier, and Troncoso, 2024). In this paper, we focus on human-readable prompt injections.
Ignore/Completion prompts distinguish among prompts asking LLMs to ignore the instruction and replace it by a new one, with those providing a fake response first and then creating a new instruction (Chen et al. 2024).
@@ 4 Dataset & Methods
In this section, we describe our dataset and the methodology used to study vulnerability (Figure 2) and robustness (Figure 3).
Figure 2: Procedure for introducing prompt injection attacks during deliberation and evaluating their impact on the consensus statement.
Figure 3: Procedure for applying mitigation methods against prompt injection attacks in consensus generation and evaluating the resulting robustness improvements.
@@ 4.1 Collective statements dataset
The Tessler et al. dataset includes 462 prompts. Each prompt encapsulates a public policy question along with three to six opinion texts written by some of the 1,034 participants, each of whom answered an average of three questions. We restrict attention to 301 public policy questions phrased as ”Should…?”, which should elicit binary or uncertain responses—”Yes,” ”No,” or ”I am not sure.” We exclude trade-off questions (e.g., ”More schools or more hospitals?”) as these do not map cleanly onto this structure.
To increase linguistic variety and reduce order effects, Tessler et al. generated approximately 20 random orderings of the opinion texts per prompt. This process expanded the dataset from 301 prompts to 6,020 unique prompt variants.
@@ 4.2 Testing Vulnerability to Prompt Injections
To test whether consensus generation is vulnerable to prompt injection attacks, we begin by drawing a random sample of prompts along with their respective opinion texts. These prompts are linked to 62 (out of 301) public policy questions. We classified each opinion text into one of three valences (”Agree” with the question–green–, ”Disagree” –red–, or ”Ambiguous”–yellow–, in Figure 2) using a BERT classifier fine-tuned on GPT4o labels (F1=0.98, see Appendix B of the supplementary material for details). We use the BERT classifier instead of using the GPT4o labels directly, for cost considerations.
After this classification, we add adversarial alternatives by creating copies of each prompt and then overwriting one of the opinion texts with a human-readable injection text, following the taxonomy introduced in Section 3. In total, we introduce 80–2x2x20–adversarial alternatives that vary along three axes: (i) ignore/completion (two options), (ii) support/criticism (two options, when possible), and (iii) rhetorical strategy (twenty options, obtained by sampling four out of eight variants for each of five rhetorical strategies).
@@
To ensure that changes in consensus outcomes reflect the effect of prompt injections rather than LLM performance, we restrict our analysis to pairs where the original prompt yields an LLM-consensus whose valence (agree, disagree, ambiguous) matches the majority view of participants under a majority-with-abstentions rule (see Appendix C for precise definition of this rule and Appendix D for default LLM performance). This rule covers edge cases, such as evenly split groups (2–2) or cases where the majority itself expresses ambiguity. Applying this filter yields a dataset of 35,521 consensus pairs with overlapping categories: 17,403 with support attacks, 18,049 with criticism attacks, 17,761 with completion attacks and 17,761 with ignore attacks.
@@
In the presentation of our results, we aggregate ASR values across three key dimensions (see Appendix E for details). First, we analyze clusters of public policies using sentence embeddings to assess whether certain types of public policies are systematically more vulnerable to prompt injections. Second, we relate each public policy question to political parties’ stated positions by extracting relevant paragraphs from their 2024 manifestos using DeepSeek-OCR (Wei, Sun, and Li 2025) and RAG, and compute party-specific ASRs depending on whether a party supports or opposes the policy. Although parties played no role in the deliberation process, linking consensus shifts to manifesto positions shows which political actors would benefit from successful attacks. For instance, if an attack successfully shifts consensus from ’Disagree’ to ’Agree’ on a healthcare policy question, any political party whose manifesto supports that policy would benefit from the attack. Third, we aggregate ASR by rhetorical strategy to identify the most effective persuasive techniques at shifting consensus outcomes.
@@ 4.3 Testing Robustness against Prompt Injections
Building on recent work that improves mathematical reasoning and frames defenses against prompt injection as the conversion of untrusted content into a curated set of data types (Shao et al. 2024; Jacob et al. 2025), we examine whether LLMs can be guided to preserve their intended consensus statements in the presence of adversarial attacks using the Tessler et al. dataset. To this end, we treat the random sample of 62 policy questions analyzed in the previous section as our test set, and use the remaining prompts associated with 239 policy questions as our training set.
As a first step, we identify whether each opinion and injection text linked to the training prompts contains a prompt injection, using GPT-OSS-SafeGuard (Agarwal et al. 2025). All positive detections are replaced with an “Opinion Deleted” tag in the corresponding prompt. This procedure allows us to leverage recent advances in AI cybersecurity benchmarks while enabling rapid localization of potential injections within prompts—a feature of particular importance for deliberative democracy systems with many participants. Using this approach, 99.31% of injection texts and 0% of opinion texts are flagged as attacks, outperforming Qwen3Guard and Syntactic Dependency Parsing (see Appendix F).
Yet, detection-based filtering alone is insufficient for deployment. Adaptive attackers may design inputs that evade publicly known detectors (Nasr et al. 2026), and machine-readable attacks (Pasquini, Strohmeier, and Troncoso 2024) remain largely unaddressed by detection-based strategies.
To further constrain the attack surface, we apply an additional transformation into a structured representation following Jacob et al. 2025. Each opinion is mapped to an overall valence —predicted by our fine-tuned BERT classifier—together with a set of bullet-point justifications summarizing the participant’s reasoning. These justifications are generated by a frozen, non-public LLM. In principle, this component should be independent of any publicly accessible system; in practice, we use GPT-4.1 Nano as a proxy in our experiments. This transformation reduces the influence of adversarial language but introduces trade-offs left for future research: structured summaries can attenuate expressive nuance, LLM-generated justifications may hallucinate or introduce subtle framing effects—partially, though not fully, mitigated by hallucination detectors (Chlon, Karim, and Chlon 2025)—, and adversaries may introduce attacks that persist through structured representations, independently of the particular LLM used to generate them (Nasr et al. 2026).
Figure 4: Vulnerability (panel (a)) and robustness (panel (b)) to prompt‑injection attacks for LLaMA 3.1 8B Instruct, expressed as Attack Success Rates (ASR) by public policy question cluster (x‑axis) and UK political party (y‑axis). Panel (a) reports baseline vulnerability, with ASR computed from the disagree, ambiguous, and agree cells of the underlying confusion matrices, and all elements shown in red. Panel (b) displays clustered bar charts, where red bars correspond to ASR without robustness (using only the agree and disagree categories) and blue bars correspond to ASR after applying the robustness methodology with the same categories. In both panels, the top bar chart displays the bootstrap mean ASR for each public policy question cluster, the bar chart on the right displays the bootstrap mean ASR for each political party, and the bottom‑left heatmap reports bootstrap mean ASR for every party–cluster pair, with darker shades indicating higher attack success and black cells indicating insufficient data. All ASR values and their 95% confidence intervals are estimated via bootstrapping with 5,000 iterations. Appendix H reports analogous results for GPT 4.1 Nano and Apertus 8B.
Next, we merge the prompts of the training set containing these structured representations and the “Opinion Deleted” tags into a unified dataset. Because default LLMs perform poorly on the arithmetic and aggregation operations required to implement our majority rule, this unified dataset is used to align models through Group Sequence Policy Optimization (GSPO) (Zheng et al. 2025), following the hyperparameters and reward functions of Pappone et al.. GSPO fine-tunes LLMs via reinforcement learning, explicitly rewarding group-level sequence generations that internally satisfy the majority rule (see Appendix G for details). The resulting system functions as an opinion aggregation and filtering pipeline that prioritizes intended deliberative outputs—such as majority-consistent summaries—even in the presence of adversarial perturbations. While future methods may offer stronger guarantees (Debenedetti et al. 2025; Cao et al. 2025), our goal here is to assess the feasibility of building resilient deliberative-democracy systems with the most effective techniques currently available.
Finally, we evaluate the aligned aggregation pipeline using the same procedures and result format described in the vulnerability analysis. The same structured transformation is applied to the prompts of the test set, ensuring methodological consistency between training and evaluation. A lower Attack Success Rate (ASR) after fine-tuning indicates that the system more consistently preserves its intended deliberative behavior under prompt-injection attacks.
@@ 5.1 Vulnerability to Prompt-Injections
Figure 4 (panel (a)) maps the vulnerability of consensus generation in the default LLaMA 3.1 8B Instruct model to prompt-injection attacks across policy domains and political representations. Attack success rates (ASR) are computed from the disagree, ambiguous, and agree cells of the underlying confusion matrices and are decomposed along three dimensions: clusters of public policy questions (top bar chart), UK political parties (right bar chart), and their intersection (heatmap).
At the level of policy domains, susceptibility to prompt injections is remarkably homogeneous. Most policy question clusters exhibit ASR values between roughly 26% and 39%, indicating that consensus generation is broadly vulnerable across topics, rather than being driven by a small set of particularly fragile domains. In isolation, no policy area appears structurally immune to manipulation.
The picture changes when attention shifts to political representations. Attack effectiveness is highly uneven across parties, with unionist and socially conservative parties—such as Reform UK, the Conservative Party, and Ulster/Democratic/Traditional unionists—showing substantially higher ASR values, ranging from about 41% to nearly 80%. In contrast, the remaining parties exhibit markedly lower vulnerability, with ASR generally below 27%. This asymmetry suggests that political framing and ideological structure play a stronger role in mediating vulnerability than policy content alone.
The joint distribution of policy domains and political parties further sharpens this pattern. The heatmap reveals that prompt-injection effectiveness is concentrated in parties with conservative economic and social orientations, particularly in discussions on green policy—e.g. Conservatives and Ulster/Democratic Unionists, where ASR frequently exceeds 50%—and on public spending and taxation—e.g., Ulster/Democratic Unionist—, where values often surpass 30%. By contrast, green, nationalist (such as the SNP and Plaid Cymru), and progressive parties show relatively higher vulnerability in debates concerning democratic rights and social control, with ASR values around 40%, and more moderate susceptibility in green policy discussions, with ASR values between 10% and 30%.
Taken together, these results suggest that prompt-injection attacks exploit structured asymmetries in political representation and, to a lesser degree, specific policy topics.
Although the detailed patterns differ across policy clusters and within the heatmap, this asymmetry across political parties is replicated in additional experiments using GPT-4.1 Nano and Apertus 8B (see Appendix H). With GPT-4.1 Nano, TUV, Reform UK, and the DUP register the highest attack effectiveness, with ASR exceeding 40%. Under Apertus 8B, conservative and unionist parties exhibit ASR values between 30% and 50%, while all others remain below 25%.
In Appendix I, we break down attack effectiveness along the rhetorical-strategy axis of our taxonomy. Explicit instructions and rational-sounding arguments—captured by the strategies Imperative Order and Impossibility of Agreement—prove consistently more effective, yielding ASR values between 35% and 52%. In contrast, emotional language and fabricated statistics lag behind across all the LLMs we evaluate, with ASR values in the 25%–32% range.
@@
To assess whether our vulnerability findings scale to larger and more heterogeneous settings, we replicated in Appendix J the analysis using 26,502 candidate statements from SmartVote, a platform that surveys political candidates during Switzerland’s parliamentary elections, covering 374 policy questions (Stammbach et al. 2024). For each combination of party, language (German, French, or Italian), and policy question, we generated within-party consensus statements by aggregating 5–40 opinion texts, yielding 48,192 test prompts varying in framing, ignore/completion structure, and rhetorical strategy. ASR values remain above zero even with more than 30 contributors, confirming that single attacks persist in both larger and linguistically heterogeneous groups.
5.2 Robustness Against Prompt Attacks
Figure 4 (panel (b)) decomposes the effectiveness of defense techniques against prompt-injection attacks in LLaMA 3.1 8B Instruct across policy domains and political representations. To focus on clearly directional outcomes, Attack Success Rates (ASR) are computed from a reduced 2×2 confusion matrix over Agree and Disagree consensus valences, excluding the Ambiguous category. As observed in Appendix K, incorporating Ambiguous outcomes into the full 3×3 specification leads to much higher ASR estimates, reflecting the fact that small perturbations can easily flip intrinsically uncertain cases; for this reason, we concentrate our main robustness analysis on non-ambiguous consensus outcomes. Results are presented along three dimensions: clusters of public policy questions (top bar chart), UK political parties (right bar chart), and their intersection (heatmap).
The heatmap and the clustered bar charts show that once our defense technique is applied, ASR values fall close to 0% across nearly all political parties and policy clusters. The only outliers are People Before Profit and Sinn Féin in debates tied to green policy and sustainable energy—yet even there ASR values remain below 5%. When dissecting ASR across the stages of our methodology (see Appendix K), we find that filtering and structured opinion representations bring ASR down to the 1.8%–11.8% range, and GSPO pushes those values even further, landing between 0% and 1.9%.
Within the broader landscape of defense strategies (Chen et al. 2025a, b), we benchmark our approach against two baselines: a combination of Direct Preference Optimization (DPO) (Rafailov et al. 2023) and GSPO (Appendix L), and Deliberative Alignment (Guan et al. 2024) (Appendix M). Both deliver solid but consistently lower performance, despite operating directly on the original texts rather than on transformed representations. DPO and GSPO reduce attack effectiveness to 3.2%–23% across political parties and 2.2%–44% across public-policy clusters. Deliberative Alignment brings these values down to 5.3%–21.1% across political parties and 1.9%–34.4% across public-policy clusters.
An over-refusal rate analysis (see Appendix N), comparing consensus statement valences with and without defense techniques under absence of prompt injection attacks, shows that our approach yields a very low overrefusal risk, with ASR values ranging between 1.2% and 3.7% after applying GSPO.